Martina leaves home at 9 a.m., bicycling at a rate of 24 mi/h. Two hours later, John leaves, driving at the rate of 48 mi/h. At what time will John catch up with Martina?
Answer by checkley75(3666)

24x=48(x-2)
24x=48x-96
24x-48x=-96
-24x=-96
x=-96/-24
x=4
proof
24*4=48(4-2)
96=48*2
96=96